OpenStudy (anonymous):

The figures below show two different ways of arranging four identical triangles of grey poster board on top of a white square. The square has sides equal to x + y, while the hypotenuse of each triangle is represented by the variable c.

OpenStudy (anonymous):

Hazel wrote the following statements to prove that c2 = x2 + y2. 1. Area of the four grey triangles inside figure A = 2. Area of the white square inside figure A = c2 3. Area of figure A = c2 + 2xy 4. Area of the four grey triangles inside figure B = 4xy 5. Area of the two white squares inside figure B = x2+ y2 6. Area of figure B = x2+ y2 + 4xy 7. Area of figure A = area of Figure B, hence c2 + 2xy = x2+ y2 + 4xy 8. Therefore, c2 = x2+ y2 Which is the first incorrect statement in Hazel’s proof?

OpenStudy (anonymous):

Which is the first incorrect statement in Hazel’s proof? Statement 6 Statement 7 Statement 4 Statement 5

OpenStudy (anonymous):

The first time I picked statement 5 and I was wrong

OpenStudy (amistre64):

isnt the area of the 4 grey triangles in B, xy + xy = 2xy? not 4xy ....

OpenStudy (anonymous):

I think your right, so statement 6 would be the right answer?

OpenStudy (amistre64):

1. Area of the four grey triangles inside figure A = 2. Area of the white square inside figure A = c2 3. Area of figure A = c2 + 2xy 4. Area of the four grey triangles inside figure B = 4xy <--- ????

OpenStudy (anonymous):

Yeah it would be 2xy and he wrote 4xy so that statement is wrong

OpenStudy (amistre64):

there are others that are wrong, but that seems to be the first to me

OpenStudy (anonymous):

Did you see the image of the equation btw? It was suppose to be before the statements in the actual question

OpenStudy (amistre64):

i did, but this site is still acting up ....

OpenStudy (anonymous):

whats wrong with OS?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Does the first statement mean that the 4 gray triangles inside square A are congruent ?

OpenStudy (anonymous):

first statement = the equation I posted as a picture

OpenStudy (mertsj):

4 is wrong. It says that the area of the 4 gray triangles in Figure B is 4 xy. But look at it. Each triangle has base x and height y so the area of each triangle is: \[A=\frac{1}{2}xy\] 4 triangles is certainly 4 times one triangle so: \[A(4 \Delta's)=4\times \frac{1}{2}xy=2xy\] So statement 4 is false.

OpenStudy (anonymous):

ok so statement 4 would of been the right answer?

OpenStudy (anonymous):

alright thanks merts! I'll tag u in the next one I need to review.

OpenStudy (anonymous):

can't close my question >.<